Is the following solved correctly?
-2t(t-5tsquared)+(-5tcubed) = -2tsquared+10tcubed+(-5tcubed) = -2tsquared+5tcubed

Answer by AnlytcPhil(1810)

-2t(t-5tsquared)+(-5tcubed) = -2tsquared+10tcubed+(-5tcubed) =
 -2tsquared+5tcubed

-2t(t-5t²)+(-5t³) = -2t²+10t³+(-5t³) = -2t²+5t³

Yes, that's correct.

To make the exponent ², turn on Num Lock.  Hold down Alt with your
left hand while typing 253 on the keypad at the far right of your
keyboard.  (It will not work if you type the number on the number
keys above the letter keys. You MUST type the number on the keypad.)
When you release the Alt key, the exponent will "magically"
appear on the screen at the cursor. 

To make the exponent ³, do the same except type 0179 instead of 253.

Edwin